OpenStudy (anonymous):

Solve the inequality 17 less than or equal to 3 - 2x less than or equal to 21

OpenStudy (anonymous):

\[17\le 3-2x \le21\] \[14\le -2x \le18\] \[-7\le x \le-9\]

OpenStudy (unklerhaukus):

when dividing by a negative do not forget to switch the inequality symbols around, Roachie has −7≤x≤−9 but this cannot be true as -9 is less that -7 . switching the symbols −7≥x≥−9

OpenStudy (anonymous):

Ohhh okay, thank you both.